A hollow sphere of inner radius 8.65 cm and outer radius 9.61 cm floats half-submerged in a liquid of density 871.00 kg/m3. (a) What is the mass of the sphere? (b) Calculate the density of the material of which the sphere is made
Answer #1

a) let the mass is m

for the floating object

weight of object = weight of liquid displaced by the object

m * g = 0.50 * 4pi/3 * 0.0961^3 * 871 * g

m = 0.50 * 4pi/3 * 0.0961^3 * 871

solving for m

m = 1.62 Kg

the mass of sphere is 1.62 Kg

b)

density of material = mass/volume of material

density of material = 1.62/(4pi/3 *(0.0961^3 - 0.0865^3))

density of material = 1609 Kg/m^3

the density of material is 1609 Kg/m^3